# Tensorized NeuroEvolution of Augmenting Topologies for GPU Acceleration

### v1

@@ Abstract.
doi: 10.1145/nnnnnnn.nnnnnnnisbn: 978-x-xxxx-xxxx-x/YY/MMconference: The Genetic and Evolutionary Computation Conference 2024; July 14–18, 2024; Melbourne, Australiajournalyear: 2024isbn: 978-1-4503-XXXX-X/18/06ccs: Theory of computation Evolutionary algorithmsccs: Theory of computation Vector / streaming algorithms
@@ 1. Introduction
During the past years, GPU acceleration has been a driving force behind the rapid advancements in AI, especially in the field of deep learning. The advancement in such hardware acceleration has facilitated the expansion of deep learning, with modern large language models incorporating hundreds of billions of parameters (Brown et al., 2020). Using GPUs for tasks like inference and back-propagation has led to significant speed improvements. Meanwhile, within the field of neuroevolution, there is also a growing interest in leveraging GPUs for hardware acceleration. Building upon the JAX framework (Frostig et al., 2018), some pioneering works such as EvoJAX (Tang et al., 2022), evosax (Lange, 2023), and EvoX (Huang et al., 2023) represent this trend, seeking to tap into the robust computational capabilities of GPU to reduce the runtime of neuroevolution algorithms, especially when handling large population sizes or extensive problem scales.
@@ 2.1. NeuroEvolution of Augmenting Topologies
  • Species-based Population Segmentation: NEAT categorizes its entire population into individual species. Conventional genetic procedures, such as selection, mutation, and crossover, are executed independently within each species. This approach serves dual purposes: it not only shields fledgling and potentially advantageous network structures from premature extinction, but also amplifies the diversity of solutions within the evolutionary exploration.
@@ 3.1. Tensorized Encodings
where $|\cdot|$ is the cardinality of a set, $\texttt{len}(\cdot)$ is the length of a one-dimensional tensor, and $\bm{n}_{i}$ and $\bm{c}_{j}$ are the tensor representations of the $i$-th node and $j$-th connection, respectively.
@@ 3.2.1. Node Modification
where $r_{i}$ denotes the index of the first NaN row in $\hat{\bm{N}}$, $\bm{n}_{\text{new}}$ stands for the tensor representation of the node being added, $\cdot$] represents tensor slicing, and $\leftarrow$ indicates the assignment operation.
@@ 3.2.3. Attribute Modification
NEAT is designed not only to modify the network structures but also the internal attributes, either in a node or a connection. Specifically, given a node $n=(k,\text{attr}_{1},\text{attr}_{2},\ldots)$, where $k\in\mathbb{N}$ is the historical marking, and $\text{attr}_{i}$ is the $i$-th attribute of $n$, when modifying the $j$-th attribute in a node, the transformation can be represented as:
@@ 4.1. JAX-based Hardware Acceleration
JAX (Frostig et al., 2018) is an open-source numerical computing library, offering APIs similar to NumPy (Harris et al., 2020) and enabling efficient execution across various hardware platforms (CPU/GPU/TPU). Leveraging the capabilities of XLA, JAX facilitates the transformation of numerical code into optimized machine instructions. The optimization techniques provided by JAX have supported the development of numerous projects in evolutionary computation, as evidenced by (Tang et al., 2022), (Lange, 2023), (Lim et al., 2022), and (Huang et al., 2023). These advancements significantly contribute to JAX’s growing popularity in the scientific community.
@@ 4.2. User-friendly Interfaces
Designed with user-friendly interfaces, TensorNEAT provides mechanisms for adjusting algorithms to specific requirements. It offers an extensive set of hyperparameters, allowing users to fine-tune various computational elements of the NEAT algorithm. Additionally, its modular problem templates facilitate the integration of specialized problems. TensorNEAT also offers several open interfaces. By implementing a few functions, users can define the behavior of networks within the NEAT algorithm. A notable feature is the interface supporting the evolution of advanced network architectures, including Spiking Neural Networks (Ghosh-Dastidar and Adeli, 2009) and Binary Neural Networks (Hubara et al., 2016). Details on the hyperparameters and the interfaces are elaborated in Appendix B and Appendix C.
@@ 5. Experiment
The adaptability of TensorNEAT was further validated by evaluating its performance across various hardware configurations, including the AMD EPYC 7543 CPU and a selection of mainstream GPU models. These evaluations were conducted in the Cart Pole environment, with a consistent population size of $10,000$. The total wall-clock time was recorded over $100$ generations of the algorithm, and the results are summarized in Table 1. In every GPU configuration, TensorNEAT demonstrated a significant performance advantage over NEAT-Python. Notably, with the RTX 4090, TensorNEAT achieved a speedup surpassing $500\times$ in the Halfcheetah environment. It is also noteworthy that TensorNEAT realized a speedup on the same CPU device in comparison to NEAT-Python, especially in the more complex Halfcheetah environment.
@@ Appendix A NeuroEvolution of Augmenting Topologies
0:$P$ (population size), II (number of input nodes), $O$ (number of output nodes), $f_{\text{target}}$ (target fitness value), $G$ (maximum number of generations)
0:bestbest
$pop\leftarrow$ initialize $P$ networks with II, $O$
@@
  $fit\leftarrow$ evaluate fitness values of PopPop
@@
  $species\leftarrow$ divide poppop by distances between networks
@@
  for $s$ in speciesspecies do
   $c\leftarrow$ determine the number of new individuals by fitfit

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2024; Melbourne, VIC, Australiabooktitle: Genetic and Evolutionary Computation Conference (GECCO ’24), July 14–18, 2024, Melbourne, VIC, Australiadoi: 10.1145/3638529.3654210isbn: 979-8-4007-0494-9/24/07ccs: Theory of computation Evolutionary algorithmsccs: Theory of computation Vector / streaming algorithms
@@ 1. Introduction
During the past years, GPU acceleration has been a driving force behind the rapid advancements in AI, especially in the field of deep learning. The advancement in such hardware acceleration has facilitated the expansion of deep learning, with modern large language models incorporating hundreds of billions of parameters (Brown et al., 2020). Using GPUs for tasks like inference and back-propagation has led to significant speed improvements. Meanwhile, within the field of neuroevolution, there is also a growing interest in leveraging GPUs for hardware acceleration. Building upon the JAX framework (Frostig et al., 2018), some pioneering works such as EvoJAX (Tang et al., 2022), evosax (Lange, 2023), and EvoX (Huang et al., 2024) represent this trend, seeking to tap into the robust computational capabilities of GPU to reduce the runtime of neuroevolution algorithms, especially when handling large population sizes or extensive problem scales.
@@ 2.1. NeuroEvolution of Augmenting Topologies
  • Species-based Population Segmentation: NEAT categorizes its entire population into individual species. Conventional genetic procedures, such as selection, mutation, and crossover, are executed independently within each species. This approach serves dual purposes: it not only shields potentially advantageous network structures from premature extinction, but also amplifies the diversity of solutions within the evolutionary exploration.
@@ 3.1. Tensorized Encodings
where ||$\cdot$|| ⋅ | is the cardinality of a set, $\texttt{len}(\cdot)$ is the length of a one-dimensional tensor, and $\bm{n}_{i}$ and $\bm{c}_{j}$ are the tensor representations of the $i$-th node and $j$-th connection, respectively.
@@ 3.2.1. Node Modification
where $r_{i}$ denotes the index of the first NaN row in $\hat{\bm{N}}$, $\bm{n}_{\text{new}}$ stands for the tensor representation of the node being added, []delimited-[][$\cdot$][ ⋅ ] represents tensor slicing, and $\leftarrow$ indicates the assignment operation.
@@ 3.2.3. Attribute Modification
NEAT is designed not only to modify the network structures but also the internal attributes, either in a node or a connection. Specifically, with node $n=(k,\text{attr}_{1},\text{attr}_{2},\ldots)$, where $k\in\mathbb{N}$ is the historical marking, and $\text{attr}_{i}$ is the $i$-th attribute of $n$, when modifying the $j$-th attribute in a node, the transformation can be represented as:
@@ 4.1. JAX-based Hardware Acceleration
JAX (Frostig et al., 2018) is an open-source numerical computing library, offering APIs similar to NumPy (Harris et al., 2020) and enabling efficient execution across various hardware platforms (CPU/GPU/TPU). Leveraging the capabilities of XLA, JAX facilitates the transformation of numerical code into optimized machine instructions. The optimization techniques provided by JAX have supported the development of numerous projects in evolutionary computation, as evidenced by (Tang et al., 2022), (Lange, 2023), (Lim et al., 2022), and (Huang et al., 2024). These advancements significantly contribute to JAX’s growing popularity in the scientific community.
@@ 4.2. User-friendly Interfaces
Designed with user-friendly interfaces, TensorNEAT provides mechanisms for adjusting algorithms to specific requirements. It offers an extensive set of hyperparameters, allowing users to fine-tune various computational elements of the NEAT algorithm. Additionally, its modular problem templates facilitate the integration of specialized problems. TensorNEAT also offers several open interfaces. By implementing a few functions, users can define the behavior of networks within the NEAT algorithm. A notable feature is the interface supporting the evolution of advanced network architectures, including Spiking Neural Networks (Ghosh-Dastidar and Adeli, 2009) and Binary Neural Networks (hubara2016binarized). Details on the hyperparameters and the interfaces are elaborated in Appendix B and Appendix C.
@@ 5. Experiment
The adaptability of TensorNEAT was further validated by evaluating its performance across various hardware configurations, including the AMD EPYC 7543 CPU and a selection of mainstream GPU models. These evaluations were conducted in the Cart Pole environment, with a consistent population size of $10,000$. The total wall-clock time was recorded over $100$ generations of the algorithm, and the results are summarized in Table 1. In every GPU configuration, TensorNEAT demonstrated a significant performance advantage over NEAT-Python. Notably, with the RTX 4090, TensorNEAT achieved a speedup surpassing 500×500$\times$500 × in the Halfcheetah environment. It is also noteworthy that TensorNEAT realized a speedup on the same CPU device in comparison to NEAT-Python, especially in the more complex Halfcheetah environment.
@@ Appendix A NeuroEvolution of Augmenting Topologies
0:  $P$ (population size), $I$ (number of input nodes), $O$ (number of output nodes), $f_{\text{target}}$ (target fitness value), $G$ (maximum number of generations)
0:  $best$
  $pop\leftarrow$ initialize $P$ networks with $I$, $O$
@@
     $fit\leftarrow$ evaluate fitness values of $Pop$
@@
     $species\leftarrow$ divide $pop$ by distances between networks
@@
     for $s$ in $species$ do
        $c\leftarrow$ determine the number of new individuals by $fit$